# Classifying fusion rules of anyons or SymTFTs: A general algebraic formula for domain wall problems and quantum phase transitions

### v2

@@ III Preserved sectors and unbroken subalgebra
It is worth reminding that the sum of the number of symmetry operator in $\mathbf{A^{\prime}}$, $n^{\prime}$, and that of modules $\{|\alpha\rangle\rangle\}_{\alpha\in\mathbb{S}_{\rho}}$, $n-n^{\prime}$, are equal to that of UV symmetry operators, $n=(n-n^{\prime})+n^{\prime}$. It is analogous to the Nambu-Goldstone (NG) theorem[132, 133, 134], and we note the corresponding phenomenology at the present stage. Historically, the connection between the pseudo-NG modes and Higgs modes has been studied in the high-energy physics phenomenology[135, 136, 137, 138], and we cite reviews[139, 140]. When assuming the perturbative expression of the massless flow, the ideal which corresponds to the massive part in the massless flow, can be interpreted as pseudo-Nambu-Goldstone modes, and they are gapped out gradually by adding the strength of the coupling constant of the relevant perturbations. The slow transition property of the massless flow can be found in numerical simulations, for example, in [141]. We note the algebraic dimension of the ideal or pseudo-NG modes as $n_{\text{pNG}}=n-n^{\prime}$.
In the dual massive RG, which can be obtained by changing the sign of the coupling constant of perturbations deriving the massless flow, the ideal appears as a module or degenerate ground states with spontaneous symmetry breaking. In this situation, the unbroken symmetry $\mathbf{A}_{\text{ub}}$ relates these ground states to each other, and $\mathbf{A}_{\text{ub}}$ corresponds to the NG modes. In general, the algebraic dimension of $\mathbf{A}_{\text{ub}}=\mathbf{A^{\prime}}_{\text{ub}}$ or operator couting of the NG modes is smaller than $n^{\prime}$ because of the inclusion $\mathbf{A^{\prime}}_{\text{ub}}\subset\mathbf{A^{\prime}}$. Hence, the following inequality holds,
@@
Because the detection of the unbroken symmetry and its action on the module are fundamental, the arguments by Graham-Watts [98] are important for further studies, and we note related recent papers again[92, 99, 100, 26, 32, 101, 102, 33].
IV Extension, orbifolding and isomorphisms
@@
V Conclusion and discussion
VI Acknowledgement
The author thanks Gerard Watts and Ingo Runkel for helpful discussions and for notifying the author of the elementary but fundamental method of solving the ring homomorphism by studying the representation by idempotents. The author thanks Yuma Furuta, Taishi Kawamoto, and Shinichiro Yahagi for related collaborations. The author thanks the support from NCTS.
$\epsilon\times\epsilon$ $=I,$ (19)
$\epsilon\times\sigma_{\text{Bulk}}$ $=\sigma_{\text{Bulk}},$ (20)
$\sigma_{\text{Bulk}}\times\sigma_{\text{Bulk}}$ $=I+\epsilon,$ (21)
$I^{\prime}$ $=1\otimes\left(\frac{I+\epsilon}{2}\right)$ (22)
$\psi^{\prime}$ $=\psi\otimes\left(\frac{I+\epsilon}{2}\right),$ (23)
$e^{\prime}$ $=1\otimes\frac{\sigma_{\text{Bulk}}}{\sqrt{2}},$ (24)
$m^{\prime}$ $=\psi\otimes\frac{\sigma_{\text{Bulk}}}{\sqrt{2}}$ (25)
We also note that there exist related but different connections between two copies of the chiral Ising model and chiral Majorana fermion model[184, 10]. We distinguish the two copies of the chiral Ising model by upper index $\{1\}$ and $\{2\}$ and denote the UV theories as $\{I^{i},\psi^{\{i\}},\sigma^{\{i\}}\}$. Then the following ring homomorphism holds,
$\rho(I^{\{1\}}\otimes I^{\{2\}})$ $=I^{\prime},$ (26)
$\rho(\psi^{\{1\}}\otimes I^{\{2\}})$ $=\psi^{\prime},$ (27)
$\rho(I^{\{1\}}\otimes\psi^{\{2\}})$ $=\psi^{\prime},$ (28)
$\rho(\psi^{\{1\}}\otimes\psi^{\{2\}})$ $=I^{\prime},$ (29)
$\rho(\sigma^{\{1\}}\otimes\sigma^{\{2\}})$ $=e^{\prime}+m^{\prime}.$ (30)
It should be stressed that we have taken subalegebra of the two copies of the Ising theory, excluding $\sigma^{[1]}\otimes I^{[2]}$, $I^{[1]}\otimes\sigma^{[2]}$$\sigma^{[1]}\otimes\psi^{[2]}$, $\psi^{[1]}\otimes\sigma^{[2]}$, and this exclusion is also a consequence of half-integral condition in forming the connected etale algebra. Because the IR theory does not have the algebraic dimension of the double semion model, $4$, the homomorphism is not surjective and unconventional from the context of massless RG and gapped domain wall. However, by introducing the relation $\sigma^{\prime}=\frac{e^{\prime}+m^{\prime}}{\sqrt{2}}$, the homomorphism, $\rho:\prod_{i}\mathbf{A}^{\{i\}}\rightarrow\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$ becomes surjective. This identification inevitably induces the unusual coefficient $\sqrt{2}$. In other words, the object $\mu=\frac{e^{\prime}-m^{\prime}}{\sqrt{2}}$ is emergent from this view. However, this emergence stems from the discrepancy between the bosonic UV and fermionic IR theories, and one can exclude this emergence by treating either the bosonic UV and IR theories or the fermionic UV and IR theories.
$\rho(\mu^{\{1\}}\otimes\mu^{\{2\}})=e^{\prime}-m^{\prime}$ (31)

### v3

@@ III Preserved sectors and unbroken subalgebra
It is worth reminding that the sum of the number of symmetry operator in $\mathbf{A^{\prime}}$, $n^{\prime}$, and that of modules of the dual massive RG, $\{|\alpha\rangle\rangle\}_{\alpha\in\mathbb{S}_{\rho}^{\text{c}}}$ with $n-n^{\prime}$ objects, are equal to that of UV symmetry operators, $n=(n-n^{\prime})+n^{\prime}$. It is analogous to the Nambu-Goldstone (NG) theorem[179, 180, 181], and we note the corresponding phenomenology at the present stage. Historically, the connection between the pseudo-NG modes and Higgs modes has been studied in the high-energy physics phenomenology[182, 183, 184, 185], and we cite reviews[186, 187]. We also note a review [188] for readers in condensed matter. When assuming the perturbative expression of the massless flow, the ideal $\mathbb{S}_{\rho}^{\text{c}}$ which corresponds to the massive part in the massless flow, can be interpreted as pseudo-Nambu-Goldstone modes, and they are gapped out gradually by adding the strength of the coupling constant $(\lambda>0)$ of the relevant perturbations. The slow transition property of the massless flow can be found in numerical simulations, for example, in [189]. We note the algebraic dimension of the ideal or pseudo-NG modes as $n_{\text{pNG}}=n-n^{\prime}$.
In the dual massive RG $(\lambda<0)$, which can be obtained by changing the sign of the coupling constant of perturbations deriving the massless flow, the ideal appears as a module or degenerate ground states with spontaneous symmetry breaking. In this situation, the unbroken symmetry $\mathbf{A}_{\text{ub}}$ relates these ground states to each other in the massive RG, and $\mathbf{A}_{\text{ub}}$ corresponds to the NG modes (which should be distinguished from the pseudo-NG modes in the massless RG). In general, the algebraic dimension of $\mathbf{A}_{\text{ub}}=\mathbf{A^{\prime}}_{\text{ub}}$, $\text{dim}(\mathbf{A}_{\text{ub}})$, or operator couting of the NG modes is smaller than $n^{\prime}$ because of the inclusion $\mathbf{A^{\prime}}_{\text{ub}}\subset\mathbf{A^{\prime}}$. Hence, the following inequality holds,
@@
Because the detection of the unbroken symmetry and its action on the module are fundamental, the arguments by Graham-Watts [140] are important for further studies, and we note related recent papers again[136, 141, 142, 33, 40, 143, 144, 41, 157, 158, 159, 160, 161, 162, 163].
IV Conclusion and discussion
V Acknowledgement
The author thanks Gerard Watts and Ingo Runkel for helpful discussions and for notifying the author of the elementary but fundamental method of solving the ring homomorphism by studying the representation by idempotents. The author thanks Yuma Furuta, Taishi Kawamoto, and Shinichiro Yahagi for related collaborations, and Shuma Nakashiba for the related discussions. The author thanks the support from NCTS.
$\epsilon\times\epsilon$ $=I,$ (16)
$\epsilon\times\sigma_{\text{Bulk}}$ $=\sigma_{\text{Bulk}},$ (17)
$\sigma_{\text{Bulk}}\times\sigma_{\text{Bulk}}$ $=I+\epsilon,$ (18)
$I^{\prime}$ $=1\otimes\left(\frac{I+\epsilon}{2}\right)$ (19)
$\psi^{\prime}$ $=\psi\otimes\left(\frac{I+\epsilon}{2}\right),$ (20)
$e^{\prime}$ $=1\otimes\frac{\sigma_{\text{Bulk}}}{\sqrt{2}},$ (21)
$m^{\prime}$ $=\psi\otimes\frac{\sigma_{\text{Bulk}}}{\sqrt{2}}$ (22)
We also note that there exist related but different connections between two copies of the chiral Ising model and chiral Majorana fermion model[211, 16]. We also note that the related lattice gauge theoretic formulations have been studied in [79]. We distinguish the two copies of the chiral Ising model by upper index $\{1\}$ and $\{2\}$ and denote the UV theories as $\{I^{i},\psi^{\{i\}},\sigma^{\{i\}}\}$. Then the following ring homomorphism holds,
$\rho(I^{\{1\}}\otimes I^{\{2\}})$ $=I^{\prime},$ (23)
$\rho(\psi^{\{1\}}\otimes I^{\{2\}})$ $=\psi^{\prime},$ (24)
$\rho(I^{\{1\}}\otimes\psi^{\{2\}})$ $=\psi^{\prime},$ (25)
$\rho(\psi^{\{1\}}\otimes\psi^{\{2\}})$ $=I^{\prime},$ (26)
$\rho(\sigma^{\{1\}}\otimes\sigma^{\{2\}})$ $=e^{\prime}+m^{\prime}.$ (27)
It should be stressed that we have taken subalegebra of the two copies of the Ising theory, excluding $\sigma^{[1]}\otimes I^{[2]}$, $I^{[1]}\otimes\sigma^{[2]}$$\sigma^{[1]}\otimes\psi^{[2]}$, $\psi^{[1]}\otimes\sigma^{[2]}$, and this exclusion is also a consequence of half-integral condition in forming the connected etale algebra. Because the IR theory does not have the algebraic dimension of the double semion model, $4$, the homomorphism is not surjective and unconventional from the context of massless RG and gapped domain wall. However, by introducing the relation $\sigma^{\prime}=\frac{e^{\prime}+m^{\prime}}{\sqrt{2}}$, the homomorphism, $\rho:\prod_{i}\mathbf{A}^{\{i\}}\rightarrow\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$ becomes surjective. This identification inevitably induces the unusual coefficient $\sqrt{2}$  [^4] Moreover, this number corresponds to the quantum dimension of $\sigma^{\prime}$. We study related phenomena elsewhere.. In other words, the object $\mu=\frac{e^{\prime}-m^{\prime}}{\sqrt{2}}$ is emergent from this view. However, this emergence stems from the discrepancy between the bosonic UV and fermionic IR theories, and one can exclude this emergence by treating either the bosonic UV and IR theories or the fermionic UV and IR theories.
$\rho(\mu^{\{1\}}\otimes\mu^{\{2\}})=e^{\prime}-m^{\prime}$ (28)
Appendix B Extension, orbifolding and isomorphisms